OpenStudy (anonymous):

If b^p=N, then which of the following statements is also true? A.) logbN=p B.) logNb=p C.) logpN=b D.) N 1/b=p

OpenStudy (anonymous):

2.) Simplify the expression: (a^2 b^5 c^3)^4 ( a^3 b^3)^3 (a^2 c)^2 A.) a^13 b^19 c^13 B.) 6(a^8 b^12 c^4) C.) a^21 b^24 c^10) D.) a^21 b^29 c^14

OpenStudy (math&ing001):

Did you know that \(\log(a^p) = p*log(a)\) and that \(\log _{a}(a)=1\)

OpenStudy (anonymous):

I didn't

OpenStudy (math&ing001):

Well you can use it now to answer question n°1 :)

OpenStudy (anonymous):

Thank you!

OpenStudy (math&ing001):

Welcome ! As for question n°2, we first distribute then we regroup like terms. For distribution remember that \((a^p)^q=a^{p.q}\), and when regrouping we use \(a^p*a^q=a^{p+q}\).

OpenStudy (math&ing001):

Btw, what did you get for n°1 ?

OpenStudy (anonymous):

#1 is A?

OpenStudy (math&ing001):

Correct !

OpenStudy (anonymous):

I don't really get #2 still

OpenStudy (math&ing001):

Well here's a simplified example: (a^2 b^3)^4 (a^3 b^4)^2 = a^(2*4) b^(3*4) a^(3*2) b^(4*2) = a^8 b^12 a^6 b^8 = a^(8+6) b^(12+8) = a^14 b^20

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

so it's b?

OpenStudy (math&ing001):

Can you show me the steps like I did ?

OpenStudy (math&ing001):

Btw it's not B.